OpenStudy (anonymous):

x+9=5(4x-2) WILL GIVE MEDAL FOR CORRECT ANSWER!

OpenStudy (anonymous):

x+9=20x-2 21x=7 x=3

OpenStudy (anonymous):

i agree

OpenStudy (anonymous):

wiseolmoosey

OpenStudy (anonymous):

would you mind helping me on a question as well?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

Wayne is buying a house for $205,000. He is financing $175,000 and obtained a 30-year, fixed-rate mortgage with a 6.425% interest rate. How much are his monthly payments?

OpenStudy (anonymous):

Answer choices are : A.11/19,B.-1, C.1, D. -1/19

OpenStudy (anonymous):

how is that possible?

OpenStudy (anonymous):

oh wait johnnyt your wrong its x+9=20x-10 19=19x 1=x

OpenStudy (anonymous):

haha miss multiplication

OpenStudy (anonymous):

It's C

OpenStudy (anonymous):

oops forgot to distribute to both

OpenStudy (anonymous):

my bad

OpenStudy (anonymous):

no probs but about your question

OpenStudy (anonymous):

Wayne is buying a house for $205,000. He is financing $175,000 and obtained a 30-year, fixed-rate mortgage with a 6.425% interest rate. How much are his monthly payments? Sadly I have no idea this reminds me of the my last lessons for Algebra 2 Honors the Sum and stuff

OpenStudy (anonymous):

lol its alright, (:

OpenStudy (anonymous):

thanks anyways though

OpenStudy (anonymous):

no problem anytime!

OpenStudy (anonymous):

Well not to not get the asnwer anytime but to help anytime :P